Have a review to do for an upcoming text and I forgot how to solve:
Solve 3^(x^(2)+4x)) = 1 / 27

Answer by Alan3354(69443)

Solve 3^(x^(2)+4x)) = 1 / 27
------------
1/27 = 3^-3 so
x^2 + 4x = 3
x^2 + 4x + 3 = 0
------
x = -1
x = -3